# Transferable Mask Transformer: Cross-domain Semantic Segmentation with Region-adaptive Transferability Estimation

### v1

@@ 3.2 Measuring Transferability with Proxy A-distance
=$\mathcal{L}=\{(\bm{x}_{\mathrm{S}}^{i},0)\}_{i=1}^{n}\cup\{(\bm{x}_{\mathrm{T}}^{i},1)\}_{i=1}^{n^{\prime}},$ (4)
@@ 4.1 Adaptive Cluster Transferability Estimator
[…]
@@
[…]
@@ 4.2 Transferability Guided Masked Attention
Linear Full LEEP OTCE Ours Linear Full LEEP OTCE Ours Linear Full LEEP OTCE Ours
  C→B 55.0 62.2 62.8 63.0 64.5 84.1 88.0 88.5 88.8 90.2 68.4 73.8 74.2 74.5 75.8
C→M 67.3 70.7 71.1 71.4 72.7 87.9 90.3 90.8 91.1 92.5 79.9 80.7 81.1 81.4 82.8
C→S 43.8 75.8 76.2 76.5 77.8 74.8 90.9 91.3 91.7 93.2 64.1 85.2 85.5 85.8 87.1
C→G 56.7 69.2 69.5 69.8 71.2 75.4 86.2 86.8 87.0 88.5 70.4 79.9 80.2 80.5 81.7
  B→C 47.3 54.7 55.2 55.5 57.5 84.9 87.2 87.8 88.0 90.2 58.4 67.5 67.9 68.2 70.2
B→M 47.0 46.9 47.3 47.6 49.0 83.4 84.6 85.0 85.3 87.5 58.1 57.1 57.5 57.8 59.8
B→S 29.3 53.0 53.4 53.7 55.8 59.6 82.0 82.4 82.7 84.5 36.6 63.4 63.8 64.0 66.4
B→G 39.8 52.0 52.5 52.8 54.8 70.9 81.9 82.3 82.6 84.3 50.7 61.8 62.2 62.5 64.3
  M→C 72.4 73.6 73.2 72.8 73.0 91.7 92.3 92.0 91.3 91.9 83.9 84.9 84.6 84.3 84.5
M→B 61.9 63.0 62.7 62.3 62.5 88.1 88.9 88.6 88.3 88.5 73.5 73.1 72.8 72.5 72.7
M→S 48.8 67.3 68.5 69.8 F 69.5 79.9 91.0 92.3 93.5 93.2 68.2 84.9 85.9 87.3 87.0
M→G 64.4 71.7 72.5 73.8 73.5 82.6 89.0 90.1 91.5 91.7 78.1 81.7 82.5 83.9 83.7
  S→C 31.1 50.9 51.2 51.5 53.5 73.8 85.8 86.1 86.3 88.0 38.9 62.5 62.9 63.2 65.7
S→B 22.7 37.6 38.0 38.2 41.1 65.1 80.2 80.3 80.4 82.3 28.3 46.0 46.4 46.7 48.5
S→M 28.7 43.7 44.0 44.3 44.0 73.4 82.3 82.6 82.8 84.5 36.2 54.0 54.3 54.6 56.1
S→G 28.4 49.7 50.0 50.3 51.8 63.2 78.8 79.2 79.6 81.4 39.5 60.2 60.5 60.8 62.2
  G→C 38.3 57.0 57.5 57.8 59.5 77.5 87.4 87.7 87.9 89.5 47.6 68.9 69.2 69.5 71.0
G→B 30.2 42.3 42.8 43.0 44.8 72.5 82.1 82.5 82.8 84.2 36.5 51.2 51.7 52.0 54.0
G→M 40.9 48.4 48.8 49.0 51.0 80.3 84.9 85.2 85.5 86.8 51.5 59.7 60.0 60.3 62.0
G→S 27.8 57.2 57.6 57.9 59.5 62.9 84.3 84.6 84.9 86.3 40.1 67.7 68.0 68.1 69.8
  Avg 44.1 57.3 57.7 58.0 59.4 76.6 85.9 86.3 86.6 88.0 55.4 68.2 68.6 68.9 70.3

### v2

@@ 3.2 Measuring Transferability with Proxy A-distance
$\mathcal{L}=\{(\bm{x}_{\mathrm{S}}^{i},0)\}_{i=1}^{n}\cup\{(\bm{x}_{\mathrm{T}}^{i},1)\}_{i=1}^{n^{\prime}},$ (4)
@@ 4.1 Adaptive Cluster Transferability Estimator
Adaptive Clustering Method.  To capture inter-regional structural differences while preserving intra-regional structural and semantic consistency, our approach generates adaptively sized regions that align closely with image boundaries. As illustrated in Fig.3, we propose an adaptive clustering method inspired by the classic works on superpixel learning [1, 17], which significantly reduces the time complexity by confining the computation range of the pixel similarity matrix. We perform an iterative grouping algorithm that takes a feature map $\bm{\chi}$ as input, and gives out a region-level feature map $\bm{f}$, where $R$ denotes the set of regions. We define a soft assignment $\text{\large$a$}\colon P\times R\to[0,1]$, where $\text{\large$a$}(p,r)$ indicates the probability of pixel $p$ being assigned to region $r$.
@@
$\text{\large$a$}((x,\!y),(i,\!j))=\mathrm{I}[\text{$(i\!-\!1)c\hskip-0.83328pt<\hskip-0.83328ptx\hskip-0.83328pt\leqslant\hskip-0.83328ptic\text{ and }(j\!-\!1)r\hskip-0.83328pt<\hskip-0.83328pty\hskip-0.83328pt\leqslant\hskip-0.83328ptjr$}].$ (6)
@@ 4.2 Transferability Guided Masked Attention
[…]